OpenStudy (anonymous):

Express answers in simplest exact form. A 10" diameter pumpkin pie is cut into six equal servings. What is the area of the top of each piece of pie? A = ? PI inches

OpenStudy (across):

Do you know what the formula for the area of a circle's sector is?

OpenStudy (anonymous):

Yes

OpenStudy (across):

\[A_s=\frac{\theta}{360}\pi r^2\]

OpenStudy (across):

The pie is cut into 6 equal servings. Therefore, each sector has an angle of 360/6=60.

OpenStudy (anonymous):

so whats my answer sorry i dont have my Calculator or pin

OpenStudy (across):

\[A_s=\frac{60}{360}\pi 5^2\]\[A_s=\frac{1}{6}\pi 25\]\[A_s=\frac{25}{6}\pi\]

OpenStudy (across):

No need of a calculator...

OpenStudy (anonymous):

o ok